# Learning to Control Unknown Strongly Monotone Games

### v1

@@ Abstract
Consider $N$ players each with a $d$-dimensional action set. Each of the players’ utility functions includes their reward function and a linear term for each dimension, with coefficients that are controlled by the manager. We assume that the game is strongly monotone, so if each player runs gradient descent, the dynamics converge to a unique Nash equilibrium (NE). The NE is typically inefficient in terms of global performance. The resulting global performance of the system can be improved by imposing $K$-dimensional linear constraints on the NE. We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE. However, this requires knowing the players’ reward functions and their action sets. Obtaining this game structure information is infeasible in a large-scale network and violates the users’ privacy. To overcome this, we propose a simple algorithm that learns to shift the NE of the game to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm, which is based on two time-scale stochastic approximation, guarantees convergence with probability 1 to the set of NE that meet target linear constraints. We then provide a mean square convergence rate of $O(t^{-1/4})$ for our algorithm. This is the first such bound for two time-scale stochastic approximation where the slower time-scale is a fixed point iteration with a non-expansive mapping. We demonstrate how our scheme can be applied to optimizing a global quadratic cost at NE and load balancing in resource allocation games. We provide simulations of our algorithm for these scenarios.
@@ 1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions. Agents do not see the bigger picture therefore optimize their local objectives instead. The result is often an inefficient equilibrium. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy. As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5, 6, 7, 8, 9].
Game theory is a natural way to tackle distributed decision-making [10, 11, 12]. The game structure formalizes the effect of the interaction between users. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they have no information regarding the reward functions of their peers. A Nash equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [13, 14, 15], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
@@
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and also maintains user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We first prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of $O(t^{-1/4})$ for the linear constraint. Due to the convex and compact action sets of the players, our slower time-scale is a fixed point iteration with a non-expansive mapping as opposed to the stronger contractive mapping assumed to analyze convergence rates in the stochastic approximation literature [16, 17]. Our bound is the first of this kind for a two time-scale stochastic approximation scheme with a non-expansive mapping. Apart from tools from stochastic approximation, we employ variational inequalities and the Krasnosel’skiĭ–Mann algorithm in our analysis. Therefore, our simple algorithm offers a theoretically principled game control scheme for strongly monotone games.
@@ 1.1 Related Work
Distributed optimization [18, 19, 20] is concerned with optimizing a target objective distributedly over a network of agents. However, these agents are non-interacting computational units so there is no game structure between them. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible. Moreover, distributed optimization entails significant communication overhead between agents.
@@
Distributed algorithms that converge to a NE have been studied in [21, 22, 23], including in monotone games [24, 25, 26]. The players in our game use simple stochastic gradient play to converge to a NE. However, if the game is not controlled, the NE is typically inefficient from a global point of view. Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
@@
Compared to mechanism design approaches [27, 28, 29], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only receives the sum of actions (loads) of all players.
@@
Our approach resembles the idea of a Stackelberg game [30, 31, 32, 33], where the manager is the leader and the players are the followers (usually a single follower is considered). However, game-theoretic algorithms converge to a NE that can be inefficient from an optimization point of view. Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
@@
Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [34, 35]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” also introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints. Hence, the type of feedback that the manager needs is about the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
@@
The most closely related literature to our work is that on intervention and control of games [36, 37, 38, 39, 40, 41, 42, 43, 44]. In [45] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [46, 40], that considered the general approach of game-theoretic control. In [36], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [47, 48, 49], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., their geometric locations). Additionally, the pricing scheme in [47, 48, 49] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [41], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
@@
There have been works obtaining convergence rate for contractive two time-scale stochastic approximation [50, 16]. These works assume that both the time-scales are fixed point iterations with contractive mappings. We consider a more general case where the slower timescale has a non-expansive mapping.
@@
A preliminary conference version, limited to load balancing in resource allocation, was published in [51]. In this paper, we extend the framework in [51] to linear constraint in strongly monotone games. In particular, the NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$. Our convergence guarantees are with probability 1 instead of in L2 like in [51]. Last but not least, we provide a finite time convergence error bound. Our novel results require new analysis based on non-expansive mappings that has limited overlap with the analysis in [51].
1.2 Outline
@@
The rest of the paper is organized as follows. In Section 2 we formulate our game control problem and detail our assumptions. In Subsection 2.1 we detail two applications that fall under our general model. In Section 3 we describe the algorithm that the manager uses to control the game and state our main results regarding its performance guarantees. In Section 4 we break the proofs of our main results into lemmas. In Section 5 we provide numerical simulations for the application described earlier. Then, In Section 6 we conclude the paper and propose future research directions.
1.3 Notation
@@ 2 Problem Formulation
Consider a set of players $\mathcal{N}=\{1,2,\ldots,N\}$, where player $n$ chooses an action $\boldsymbol{x}_{n}=(x_{n}^{1},\ldots,x_{n}^{d})\in\mathcal{X}_{n}$. Here $\mathcal{X}_{n}\subset\mathbb{R}^{d}_{+}$ is a convex and compact set such that $\boldsymbol{0}\in\mathcal{X}_{n}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$. We use $\boldsymbol{x}\in\mathcal{X}$ to denote the concatenation of all action vectors. Then the utility for each player is defined as:
$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{d}x_{n}^{d}.$ (1)
@@
Our game evolves in discrete time, with players taking actions $\boldsymbol{x}_{n,t}$, given the controlled coefficients $\boldsymbol{\beta}_{n,t}$, at time $t$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the objective of the game manager (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., the game is uncontrolled), then the players will eventually converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
@@
$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c}.$ (5)
Here $\mathcal{F}_{t}$ is the filtration that summarizes the past, i.e.,
@@ 2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global objective of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$ which they wish to minimize. All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $A=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a local optimum.
@@ 2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running DSM to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [56] and control of parking resources [57]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources. We have $d$ constraints i.e., $K=d$ with the idea of each linear constraints being to meet the target loads for each resource $i$, i.e.,
@@
The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize for application-specific objectives such as system efficiency, regulations, and operation costs. The controlled coefficients $\boldsymbol{\beta}$ can be thought of as incentives to use a resource in this application. The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
A special case of this is resource allocation without weights, studied in [51]. In this case the linear constraints are
[…]
Generalizing to a weighted sum of actions allows modeling of cases such as power control in wireless networks and load balancing in data-centers. In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data-centers, different users send jobs of different sizes to the different servers (resources), so their actions (how many jobs to send where) affect the loads differently.
@@ 3 Game Control Algorithm
Input: Target vector $\boldsymbol{l}^{*}$.
@@
The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$. To overcome this, we propose an online learning approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$. We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$. This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
Our algorithm can be thought of as a two time-scale stochastic approximation algorithm. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
This algorithm can also be thought of as an iterative approach to solving the constrained optimization problem using Lagrange multipliers. We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$. The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied at all times as that will require communication between agents, and hence the manager uses iterative approach to solve the constrained optimization problem. Similar approaches have been used in [58, 59] for different problems.
@@
In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [60, 61, 15]. In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
@@
The manager’s algorithm in (8) is robust against strategic manipulations of a small group of players since each player has a negligible effect on $\boldsymbol{A}\boldsymbol{x}_{t-1}$, which is its only way to impact (8). It is not clear if even a large group of players can manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players. Moreover, truthful reporting is not an issue in our algorithm since the players do not report anything, and the manager cannot observe their actions. Instead, the manager only needs to measure the instantaneous constraint violation

### v2

@@ Abstract
Consider a game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager. We assume that the game is strongly monotone, so gradient play converges to a unique Nash equilibrium (NE). The NE is typically globally inefficient. The global performance at NE can be improved by imposing linear constraints on the NE. We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE. However, this requires knowing the players’ reward functions and action sets. Obtaining this game information is infeasible in a large-scale network and violates user privacy. To overcome this, we propose a simple algorithm that learns to shift the NE to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm converges with probability 1 to the set of NE that satisfy target linear constraints. We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
@@ 1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions. Agents do not see the bigger picture therefore optimize their local objectives instead. The result is often an inefficient equilibrium. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy. As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5].
Game theory is a natural way to tackle distributed decision-making [6, 7]. The game structure formalizes the effect of the interaction between users. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they do not know the reward functions of their peers. A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [8, 9, 10], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
@@
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and maintains user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraint. Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping typically assumed to analyze convergence rates [11, 12]. Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis. Therefore, we provide a theoretically principled game control scheme for strongly monotone games.
@@ 1.1 Related Work
Distributed optimization [13, 14] is concerned with optimizing a target objective distributedly over a network of agents. However, these agents are non-interacting computational units so there is no game structure between them. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible. Moreover, distributed optimization entails significant communication overhead between agents.
@@
Distributed algorithms that converge to NE have been studied in [15, 16, 17], including in monotone games [18, 19, 20]. The players in our game use simple stochastic gradient play to converge to NE. However, if the game is uncontrolled, the NE is typically globally inefficient. Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
@@
Compared to mechanism design approaches [21, 22], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only observes the constraint violation.
@@
Our approach resembles a Stackelberg game [23, 24, 25], where the manager is the leader and the players are the followers (usually a single follower is considered). However, game-theoretic algorithms converge to a NE that can be globally inefficient. Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
@@
Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [26, 27]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints. Hence, the type of feedback the manager needs concerns the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
Our problem can be viewed as a Generalized Nash Equilibrium Problem (GNEP), where players seek to converge to a NE subject to coupled constraints [28]. Efficient computation of GNE in strongly monotone games in the noiseless setting has been widely studied [29, 30, 31]. Stochastic convergence to GNEs has recently been studied under different assumptions. In [32] players share their actions with others, and in [33] the reward function is assumed to be linear. In [34], stochastic convergence to GNE is studied more generally, resulting in a centralized algorithm where the weighted empirical average of the actions is shown to converge in expectation. We provide a distributed stochastic learning scheme that converges in probability 1 to GNE with linear coupled constraints in strongly monotone games without communication between players.
@@
Our work belongs to the literature on intervention and control of games [35, 36, 37, 38, 39, 40, 41, 42, 43]. In [44] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [45, 39], that considered the general approach of game-theoretic control. In [35], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [46, 47, 48], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., geometric locations). Additionally, the pricing scheme in [46, 47, 48] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [40], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
@@
Some works obtained convergence rates for contractive two time-scale stochastic approximation [49, 11]. These works assume that both the time-scales are fixed point iterations with contractive mappings. We consider a more general case where the slower timescale has a non-expansive mapping.
@@
An earlier conference paper [50] was limited to load balancing in resource allocation. This paper’s framework is similar to that in [50], but is extended to linear constraint in strongly monotone games. The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$. Nevertheless, the results here are completely novel and independent of [50]. Our convergence guarantees are with probability 1 instead of in L2 like in [50]. Unlike [50], we provide a finite time convergence error bound. Our novel analysis is based on two time-scale stochastic approximation with non-expansive mappings.
1.2 Outline and Notation
@@
The rest of the paper is organized as follows. Section 2 formulates our game control problem and details our assumptions. Subsection 2.1 details two applications that fall under our general model. Section 3 describes the game control algorithm and states our main results regarding its performance guarantees. Section 4 breaks the proofs of our main results into lemmas. Section 5 provides numerical simulations for the application described earlier. Finally, Section 6 concludes the paper and proposes future research directions.
@@ 2 Problem Formulation
Consider a set of players $\mathcal{N}=\{1,2,\ldots,N\}$, where player $n$ chooses an action $\boldsymbol{x}_{n}=(x_{n}^{1},\ldots,x_{n}^{d})\in\mathcal{X}_{n}$. Here $\mathcal{X}_{n}\subset\mathbb{R}^{d}_{+}$ is a convex and compact set such that $\boldsymbol{0}\in\mathcal{X}_{n}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$. We use $\boldsymbol{x}\in\mathcal{X}$ to denote the concatenation of all action vectors. The utility for each player is defined as:
$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
@@
Our game evolves in discrete time, so at time $t$ players take actions $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
@@
$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c},$ (5)
for some $M_{c}>0$ and a filtration $\mathcal{F}_{t}$ that summarizes the past:
@@ 2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
@@ 2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [55] and control of parking resources [56]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources. We have $d$ linear constraints i.e., $K=d$ and each represents a target load for each resource $i$, i.e.,
@@
The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize application-specific objectives such as system efficiency, regulations, and operation costs. The controlled coefficients $\boldsymbol{\beta}$ represent incentives to use a resource. The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
A special case is resource allocation without weights, studied in [50]. Generalizing to a weighted sum of actions allows modeling of power control in wireless networks and load balancing in data centers. In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where).
@@ 3 Game Control Algorithm
Input: Target vector $\boldsymbol{\ell^{*}}$.
@@
The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$. To overcome this, we propose an online approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$. We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$. This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
Our algorithm is a two time-scale stochastic approximation algorithm. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
Essentially, our algorithm is an iterative approach to solving the constrained optimization problem using Lagrange multipliers. We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$. The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied for all $t$ without communicating, and hence the manager uses an iterative approach instead. Similar approaches have been used in [57, 58] for different problems.
@@
In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [59, 10]. In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
@@
The manager’s algorithm in (8) is robust against strategic manipulations of a small group of players since each player has a negligible effect on $\boldsymbol{A}\boldsymbol{x}_{t-1}$, which is its only way to impact (8). We conjecture that even a large group of players cannot manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players. Moreover, truthful reporting is not an issue since the players do not report anything, and the manager cannot observe their actions. Instead, the manager only needs to measure the instantaneous constraint violation.